# Discovering Concept Directions from Diffusion-based Counterfactuals via Latent Clustering

### v1

@@ 11 Data availability
This supplementary section presents the experimental setup used to extract concept directions using the encoder from the CDCT framework (hereafter referred to as the CDCT encoder). The CDCT encoder is a Variational Autoencoder (VAE) trained on a counterfactual trajectory dataset generated via a latent diffusion model with classifier guidance, as described in Step 2 (Semantic Space Disentanglement) of the CDCT framework [15]. The training objective encourages the latent space to encode distinct, classifier-relevant transformations, facilitating more interpretable and disentangled concept representations. Further architectural and training details are provided in the original work [15].
@@
In this supplementary analysis, the same pipeline described in Section 3.2 of the main manuscript is applied to compute latent difference vectors and extract concept directions. Specifically, for each training sample not predicted as the target class, a classifier-guided counterfactual is generated, and both the factual and counterfactual images are encoded using the CDCT encoder. Unlike the LDM encoder, which yields latent embeddings of shape $4\times 32\times 32$, the CDCT encoder produces compact latent representations of shape $512\times 1\times 1$. Each latent difference is flattened into a 512-dimensional vector and normalized to unit length to produce directional vectors. Spherical K-Means clustering [29] is then applied to the set of unit-norm latent differences computed per class. The number of clusters is selected based on the silhouette coefficient to capture cohesive and distinct semantic directions. The average unit direction within each resulting cluster is interpreted as a representative concept direction for the target class.
@@
Interpretation and Evaluation of Concept Directions Using the CDCT Encoder
@@
The concept directions extracted using the CDCT encoder exhibit class-specific semantic transformations aligned with known dermoscopic features. For Melanoma, the direction (Figure 3a) darkens the lesion periphery, reduces central contrast, and softens the borders. The Basal Cell Carcinoma direction (Figure 3b) reduces pigmentation and softens edge contrast, yielding a more blended lesion appearance. In the case of Actinic Keratosis (Figure 3c), the identified direction introduces a central lightened area with a scaly texture and peripheral reddish hue. The Benign Keratosis direction (Figure 3d) induces reddish central pigmentation with uneven borders and surrounding pigment variation. For Dermatofibroma, the concept direction (Figure 3e) enhances a pale central region surrounded by a subtle pink hue, features reported in literature [32]. The direction for Vascular Lesion (Figure 3f) reveals central redness accompanied and subtle irregularities in pigment distribution, resembling vascular blushes typically observed in angiomas and angiokeratomas [33] Finally, the Squamous Cell Carcinoma direction (Figure 3g) introduces a central white, keratin-like structure surrounded by inflamed reddish tissue, consistent with dermoscopic features described by Rosendahl et al. [34].

### v2

@@ Appendix E Extraction of Concept Directions Using the CDCT Encoder
This supplementary section presents the experimental setup used to extract concept directions using the encoder from the CDCT framework (hereafter referred to as the CDCT encoder). The CDCT encoder is a Variational Autoencoder (VAE) trained on a counterfactual trajectory dataset generated via a latent diffusion model with classifier guidance, as described in Step 2 (Semantic Space Disentanglement) of the CDCT framework [varshney2025generating]. The training objective encourages the latent space to encode distinct, classifier-relevant transformations, facilitating more interpretable and disentangled concept representations. Further architectural and training details are provided in the original work [varshney2025generating].
@@
In this supplementary analysis, the same pipeline described in Section 3.2 of the main manuscript is applied to compute latent difference vectors and extract concept directions. Specifically, for each training sample not predicted as the target class, a classifier-guided counterfactual is generated, and both the factual and counterfactual images are encoded using the CDCT encoder. Unlike the LDM encoder, which yields latent embeddings of shape $4\times 32\times 32$, the CDCT encoder produces compact latent representations of shape $512\times 1\times 1$. Each latent difference is flattened into a 512-dimensional vector and normalized to unit length to produce directional vectors. Spherical K-Means clustering [hornik2012spherical] is then applied to the set of unit-norm latent differences computed per class. The number of clusters is selected based on the silhouette coefficient to capture cohesive and distinct semantic directions. The average unit direction within each resulting cluster is interpreted as a representative concept direction for the target class.
@@
Appendix F Interpretation and Evaluation of Concept Directions Using the CDCT Encoder
@@
The concept directions extracted using the CDCT encoder exhibit class-specific semantic transformations aligned with known dermoscopic features. For Melanoma, the direction (Figure 3a) darkens the lesion periphery, reduces central contrast, and softens the borders. The Basal Cell Carcinoma direction (Figure 3b) reduces pigmentation and softens edge contrast, yielding a more blended lesion appearance. In the case of Actinic Keratosis (Figure 3c), the identified direction introduces a central lightened area with a scaly texture and peripheral reddish hue. The Benign Keratosis direction (Figure 3d) induces reddish central pigmentation with uneven borders and surrounding pigment variation. For Dermatofibroma, the concept direction (Figure 3e) enhances a pale central region surrounded by a subtle pink hue, features reported in literature [agero2006conventional]. The direction for Vascular Lesion (Figure 3f) reveals central redness accompanied and subtle irregularities in pigment distribution, resembling vascular blushes typically observed in angiomas and angiokeratomas [piccolo2018dermatoscopy] Finally, the Squamous Cell Carcinoma direction (Figure 3g) introduces a central white, keratin-like structure surrounded by inflamed reddish tissue, consistent with dermoscopic features described by Rosendahl et al. [rosendahl2012dermoscopy].